# Motion Matters: Motion-guided Modulation Network for Skeleton-based Micro-Action Recognition

### v3

@@ 2.1. Micro-Action Recognition
$2.$ Skeleton-based Action Recognition
@@ 3.2. Network Overview
where $t$, $v$, and cscs denote the temporal step, joint index, and channel index, respectively.
@@ 3.4. Motion-guided Feature Modulation
To capture the subtle motion patterns in the skeleton sequence, we propose a Motion-guided Feature Modulation (MFM) module that models complex motion dynamics to guide the learning of compact spatio-temporal features. As shown in Figure 2, MFM is composed of NN stacked Motion-guided Skeletal-Temporal Formers (MSTF). Inspired by Transformer-based methods (Wang and Koniusz, 2023; Do and Kim, 2024) in skeleton-based action recognition, we utilize three parallel branches—namely, the skeletal branch, motion branch, and temporal branch—to construct compact spatio-temporal representations. The input is first normalized and projected as:
@@
(12) ${\bm{X}}_{gcm}=\frac{{\bm{X}}_{b}-\mu({\bm{X}}_{tc})}{\sigma({\bm{X}}_{tc})}\cdot(1+\gamma_{s})+\beta_{s}.$
@@
(15) ${\bm{X}}_{tcm}=\frac{{\bm{X}}_{b}-\mu({\bm{X}}_{tc})}{\sigma({\bm{X}}_{tc})}\cdot(1+\gamma_{t})+\beta_{t},$
@@
where $\cdot$] denotes channel-wise concatenation, $\texttt{Aggregation}(\cdot)$ learns different weights based on the feature distributions of each branch. This strategy helps balance contributions from spatial, temporal, and motion-specific features. To further enhance network stability, the fused features ${\bm{X}}_{agg}$ pass through a projection layer and are then added via a residual connection to the original input ${\bm{X}}_{raw}$. A FeedForward network (Vaswani et al., 2017) is subsequently applied to introduce nonlinear transformations. Overall, the stacked NN-layer Motion-guided Skeletal-Temporal Former preserves deeper motion cues, thereby strengthening the final representations ${\bm{X}}_{f}$ of micro-action movements.

### v4

@@ 2.1. Micro-Action Recognition
2.2. Skeleton-based Action Recognition
@@ 3.3. Feature Embedding
where $t$, $v$, and $cs$ denote the temporal step, joint index, and channel index, respectively.
@@ 3.4. Motion-guided Feature Modulation
To capture the subtle motion patterns in the skeleton sequence, we propose a Motion-guided Feature Modulation (MFM) module that models complex motion dynamics to guide the learning of compact spatio-temporal features. As shown in Figure 2, MFM is composed of $N$ stacked Motion-guided Skeletal-Temporal Formers (MSTF). Inspired by Transformer-based methods (Wang and Koniusz, 2023; Do and Kim, 2024) in skeleton-based action recognition, we utilize three parallel branches—namely, the skeletal branch, motion branch, and temporal branch—to construct compact spatio-temporal representations. The input is first normalized and projected as:
@@
(12) ${\bm{X}}_{gcm}=\frac{{\bm{X}}_{gc}-\mu({\bm{X}}_{gc})}{\sigma({\bm{X}}_{gc})}\cdot(1+\gamma_{s})+\beta_{s}.$
@@
(15) ${\bm{X}}_{tcm}=\frac{{\bm{X}}_{tc}-\mu({\bm{X}}_{tc})}{\sigma({\bm{X}}_{tc})}\cdot(1+\gamma_{t})+\beta_{t},$
@@
where $\cdot$] denotes channel-wise concatenation, $\texttt{Aggregation}(\cdot)$ learns different weights based on the feature distributions of each branch. This strategy helps balance contributions from spatial, temporal, and motion-specific features. To further enhance network stability, the fused features ${\bm{X}}_{agg}$ pass through a projection layer and are then added via a residual connection to the original input ${\bm{X}}_{raw}$. A FeedForward network (Vaswani et al., 2017) is subsequently applied to introduce nonlinear transformations. Overall, the stacked $N$-layer Motion-guided Skeletal-Temporal Former preserves deeper motion cues, thereby strengthening the final representations ${\bm{X}}_{f}$ of micro-action movements.